OpenStudy (kweenlexis):

7th grade math algebra to i think.

OpenStudy (anonymous):

yea?

OpenStudy (bluemiku):

is there a link?

OpenStudy (alexandervonhumboldt2):

question ms?

OpenStudy (kweenlexis):

Which literal equations are equivalent to F = ma? Choose all answers that are correct. A.\[m=Fa\] B.\[m=\frac{ F}{ a }\] C. \[a=\frac{ m }{ F }\] D. \[a=\frac{ F }{ m}\] sorry i had to type out the question.

OpenStudy (anonymous):

@kweenlexis what is the question?

OpenStudy (kweenlexis):

more the oneanswer choice

OpenStudy (kweenlexis):

@atreyu6s3x6 @AlexandervonHumboldt2 @BlueMiku

OpenStudy (anonymous):

sorry i was busy but you still need help?

OpenStudy (kweenlexis):

it's ok and yes i do.

OpenStudy (anonymous):

okay so our equation is F=ma, correct?

OpenStudy (kweenlexis):

correct.

OpenStudy (anonymous):

@kweenlexis so if we isolate each letter, we can find the equivalents, so: F=ma is we isolate "a" by dividing it from each side, we find that a=F/m is equivalent to F=ma, correct?

OpenStudy (anonymous):

by dividing "m" sorry xD

OpenStudy (anonymous):

try to isolate, "m" and find the other equivalent

OpenStudy (kweenlexis):

huh? you got me lost after you changed it to m

OpenStudy (kweenlexis):

STOP TALKING TO ME @superman38 DANG you cant see i dont want to talk to you dang

OpenStudy (anonymous):

alright, so i divided each side by m to isolate a, so i got a=F/m

OpenStudy (anonymous):

@kweenlexis

OpenStudy (kweenlexis):

ok but its more than one answer choice i rally dont get this at all

OpenStudy (anonymous):

okay so we can also divide each side by "a" to isolate "m", we get m=F/a, true?

OpenStudy (kweenlexis):

ok can you explain how you get \[m=\frac{ F }{ a }\]

OpenStudy (anonymous):

thats what i jus did haha isolate m by dividing a

OpenStudy (kweenlexis):

oh i see

OpenStudy (kweenlexis):

ok so the answers are \[a=\frac{ F }{ m }\] and \[m=\frac{ F }{ a }\] but i still don t get it. but thank you

OpenStudy (anonymous):

yes